OpenStudy (anonymous):

How do you do 2/5x+3/7=1-4/7x ?

OpenStudy (anonymous):

2x/5 + 3/7 = 1 - 4x/ 7 14x + 15 = 35 -20x 24x = 20 x = 20/24 = 5/6

OpenStudy (anonymous):

cross multi[ply

OpenStudy (anonymous):

that helps u @AEB047

OpenStudy (anonymous):

The answer is\[x=\frac{10}{17} \]